OpenStudy (anonymous):

Simplify and state the domain for the expression:

OpenStudy (anonymous):

\[\frac{ 5x^3+15x^2+10x }{ 3x^3- 3x }\]

OpenStudy (anonymous):

Domain will be all real numbers except those that make the denominator 0, so solve 3x^3-3x=0

OpenStudy (anonymous):

How would you solve that? It'd end up being 6x = 0 correct? You can't divide 6 by 0

OpenStudy (anonymous):

Or would it just be 6 because that'd be 6 over 0

OpenStudy (anonymous):

You would factor it 3x(x^2-1) so 3x=0 and x^2-1=0

OpenStudy (anonymous):

Alright thanks I got it.